### v2

Abstract
Various tasks in data science are modeled utilizing the variational regularization approach, where manually selecting regularization parameters presents a challenge. The difficulty gets exacerbated when employing regularizers involving a large number of hyperparameters. To overcome this challenge, bilevel learning can be employed to learn such parameters from data. However, neither exact function values nor exact gradients with respect to the hyperparameters are attainable, necessitating methods that only rely on inexact evaluation of such quantities. State-of-the-art inexact gradient-based methods a priori select a sequence of the required accuracies and cannot identify an appropriate step size since the Lipschitz constant of the hypergradient is unknown. In this work, we propose an algorithm with backtracking line search that only relies on inexact function evaluations and hypergradients and show convergence to a stationary point. Furthermore, the proposed algorithm determines the required accuracy dynamically rather than manually selected before running it. Our numerical experiments demonstrate the efficiency and feasibility of our approach for hyperparameter estimation on a range of relevant problems in imaging and data science such as total variation and field of experts denoising and multinomial logistic regression. Particularly, the results show that the algorithm is robust to its own hyperparameters such as the initial accuracies and step size.
keywords:
Bilevel Optimization, Bilevel Learning, Hyperparameter Optimization, Backtracking Line Search, Machine Learning.
{MSCcodes}
65K10, 90C25, 90C26, 90C06, 90C31, 94A08
1 Introduction
In this work, we consider bilevel learning, where the hyperparameter learning problem is formulated as a bilevel optimization problem [15, 35] where our goal is to solve
$\min_{\theta\in\mathbb{R}^{d}}\Bigl{\{}f(\theta)$ $\coloneqq\frac{1}{m}\sum_{i=1}^{m}g_{i}(\hat{x}_{i}(\theta))+r(\theta)\Bigr{\}}$ (1a)
$s.t.\quad{\hat{x}_{i}(\theta)}$ $\coloneqq\arg\min_{x\in\mathbb{R}^{n}}h_{i}(x,\theta),\quad i=1,\dots,m.$ (1b)
The upper-level functions $g_{i}$ are a measure of quality for the solution of the lower-level problem (1b). For example, in supervised learning, they may take the form $g_{i}(x)=\|x-x^{*}_{i}\|^{2}$ where $x^{*}_{i}$ is the desired solution of the lower-level problem.
Problems of the form (1) arise in many areas of data science, for example, when the task at hand is modeled using variational regularization approaches, which are commonly used in the fields of image reconstruction, image processing and machine vision. For instance, one can learn the weights of the regularizer in regression [33], or the parameters of the regularizer [24, 14, 32] for tasks such as image denoising, deblurring, inpainting, segmentation, and single-image super-resolution. Moreover, sophisticated data-adaptive regularizers can have millions of parameters; examples include input-convex neural networks [28, 1] and convex-ridge regularizer neural networks [22]. However, currently, neither of these data-adaptive regularizers are learned by bilevel learning due to computational difficulties. The relevance of bilevel learning in variational regularization methods is not restricted to regularizers. It can also be employed to learn the parameters in analytical deep priors [16, 2], and the sampling operator in MRI [37] and seismic imaging [17]. These tasks can also involve a large number of parameters; for example, the authors in [37] demonstrated the use of more than a million parameters. Furthermore, based on the theoretical approach introduced in [8], bilevel learning can be employed to learn the optimal structure of the regularizer.
To address the bilevel learning problem (1), one can employ model-free approaches like grid search and random search [5] when dealing with a limited set of hyperparameters and a small, bounded hyperparameter space. Moving beyond the capabilities of model-free approaches, model-based approaches such as Bayesian methods [21] and derivative-free optimization [12, 10, 18] prove to be efficient in solving (1) when a small number of hyperparameters are involved. However, here we are interested in computationally scalable algorithms that can solve bilevel learning problems potentially with millions of parameters, thus, in this work, we consider gradient-based methods. In the context of the gradient-based approach, firstly, a first-order or quasi-Newton algorithm [14, 26] is employed to solve the lower-level problem (1b). Then, gradient descent is applied to the upper-level problem (1a) after computing the gradient $\nabla f(\theta)$ (also referred to as the hypergradient) with respect to parameters $\theta$.
Under certain regularity assumptions, when the problem is sufficiently smooth [18, 33], the hypergradient can be calculated by utilizing the implicit function theorem (IFT) [18] or by using automatic differentiation (AD) [27]. Another type of algorithm called piggyback [23] utilizes AD to compute the hypergradient, and they can differ in the required assumptions on the lower-level problem (1b). This approach is favorable for nonsmooth lower-level problems. The works [6, 7] provided an extensive analysis of this method, where [7] considered approximating the hypergradient under the usage of different splitting methods, and [6] studied the convergence of piggyback AD when the lower-level problem is a saddle point problem and employed primal-dual algorithms for solving it. As an application, the authors in [9] applied piggyback AD to learn discretization of total variation. Quasi-Newton algorithms [31] are used in [34] for solving the lower-level problem and utilized their approximation of the inverse of the Hessian of the lower-level objective to replace CG.
In practice, due to the usage of numerical solvers and the large-scale nature of the problems of interest, computing the exact hypergradient is not feasible [20]; hence, it should be approximated with an accuracy that leads to optimization with inexact gradients in the upper-level problem (1a). Computing an approximate hypergradient using the IFT approach requires solving a linear system and solving it up to a tolerance using conjugate gradient (CG). The first error-bound analysis of the approximate gradient computed in this approach was done in [33] and improved in [41]. Subsequently, the error bound analysis of (IFT + CG) was further extended by providing computable a priori and a posteriori bounds in [20]. On the other hand, the error-bound analysis of approximating the hypergradient using AD was studied in [27]. Furthermore, the authors in [20] proposed a unified perspective in which AD can be seen as equivalent to IFT+CG.
1.1 Challenges
Required accuracy for hypergradient approximation
Although much work has been done on different methods of finding the hypergradient and its error bound, convergence theory for hyperparameter optimization with an inexact hypergradient has often assumed the presence of an accurate approximate hypergradient. However, in practice, approximating the hypergradient to a high degree of accuracy results in prohibitively high computational costs. Moreover, empirical evidence, as demonstrated in studies such as [20], indicates that even when the accuracy is not very high or strictly increasing under a priori assumptions like summability, inexact gradient descent on the upper-level problem can still yield progress. As elevating accuracy necessitates additional computations, implementing a dynamic strategy for selecting the required accuracy, allowing it to adjust according to the demands of the optimization, opens the possibility of reducing the overall computational cost and increase the efficiency of the optimization process.
Selecting upper-level step sizes
The convergence of inexact gradient descent on the upper-level problem has primarily been studied under the assumption of a sufficiently small fixed step size [33]. However, due to the unavailability of a closed-form solution for the lower-level problem (1b), estimating the Lipschitz constant of the hypergradient and consequently employing a method with the largest possible fixed step size is unrealistic. While line search approaches such as backtracking may seem plausible to address this issue, they typically require evaluating the exact upper-level objective value, which is unattainable since only an approximate lower-level solution is available. Additionally, line search methods like the Armijo rule [31] necessitate the approximated hypergradient to be a descent direction, which is not obvious.
1.2 Existing work and contributions
The IFT+CG approach for approximating the hypergradient, given a decreasing sequence of accuracies, was initially utilized in the hyperparameter optimization approximate gradient (HOAG) algorithm [33]. In this algorithm, convergence of the inexact gradient descent in the gradient mapping was demonstrated under the conditions of having an inexact gradient with a summable error bound and a sufficiently small step size [33]. While HOAG offered improved computational efficiency compared to methods relying on highly accurate approximate hypergradients, the assumption of summability may require more lower-level computations than necessary. Moreover, as discussed regarding the challenge of finding a suitable small enough fixed step size, these assumptions could potentially limit the practical performance of the method. In addressing the challenge of determining a suitable step size for inexact gradient descent in the upper-level, the author in [33] proposed a heuristic line search in numerical experiments. However, this line search lacked any convergence guarantee. Additionally, an accelerated algorithm of this type with a restarting scheme was developed in [40].
In [39], the authors introduced an IFT-based algorithm that performs only a single step to solve the lower-level problem. The algorithm utilizes CG with high accuracy or analytic exact inversion of the Hessian of the lower-level problem.
Moreover, some recent work has developed line search methods for single-level problems where the objective function value is computed with errors and the estimated gradient is inexact [38] and potentially stochastic [4]. However, in these methods, the error is bounded but irreducible and not controllable.
An existing adaptive and inexact framework for bilevel learning, known as DFO-LS [19], operates by solving the lower-level problem up to a specified tolerance. Subsequently, leveraging an approximate lower-level solution, the framework utilizes a derivative-free algorithm [11] tailored for solving nonlinear least-squares to update the parameters in the upper-level. This method dynamically determines the necessary accuracy for solving the lower-level problem by considering the trust-region radius of the derivative-free algorithm, ensuring that the upper-level solver can progress effectively. While this approach shows promise in tasks such as determining the parameters of total variation denoising, it faces scalability challenges as the number of parameters increases [15].
In this work, we propose a verifiable backtracking line search scheme in Section 3.1.2 that relies solely on inexact function evaluations and the inexact hypergradient. This scheme guarantees sufficient decrease in the exact upper-level objective function. We also establish the required conditions for finding a valid step size using this line search method in Section 3.1.3 and prove the convergence of inexact gradient descent in the upper-level objective function when employing our inexact line search in Section 3.2.
We present a practical algorithm in Section 3 that connects all of our theoretical results, determining the required accuracy for the inexact hypergradient to be considered a descent direction and the inexact line search (sufficient decrease condition) to be held, adaptively. As a result, our algorithm provides a robust and efficient method for bilevel learning. Furthermore, as the numerical experiments in Section 4 demonstrate, our approach outperforms state-of-the-art methods such as HOAG [33] and DFO-LS [18] in solving problems like multinomial logistic regression and variational image denoising, respectively.
1.3 Notation
[…]
2 Background
For ease of notation, we will only consider one data point and the index of the data points $i$ is omitted in the following discussion. Similarly, we do not consider the regularizer $r$. Thus, our general bilevel learning problem we will study in the main part of the paper takes the form:
$\min_{\theta\in\mathbb{R}^{d}}\{f(\theta)$ $\coloneqq g(\hat{x}(\theta))\}$ (2a)
${\hat{x}(\theta)}$ $\coloneqq\arg\min_{x\in\mathbb{R}^{n}}h(x,\theta).$ (2b)
Now, we consider the following assumptions. {assumption} We make the following assumptions on the lower level loss $h$:
  • There exist $\mu(\theta),\ L(\theta)\in\mathbb{R}$, $0<\mu(\theta)\leq L(\theta)$ such that $\mu(\theta)I\preceq\nabla_{x}^{2}h(x,\theta)\preceq L(\theta)I$. Moreover, $\nabla_{x}h$ and $\nabla_{x}^{2}h$ are continuous in $\theta$.
  • $\nabla_{x}^{2}h(x,\theta)$ is $L_{H}$ Lipschitz continuous in $x$ and $\nabla_{x\theta}^{2}h(x,\theta)$ is $L_{J}$ Lipschitz in $x$ uniformly for all $\theta$.
  • $f$ is $L_{\nabla f}$-smooth and $g$ is $L_{\nabla g}$-smooth, which means $f$ and $g$ are continuously differentiable with $L_{\nabla f}$ and $L_{\nabla g}$ Lipschitz gradients, respectively. Moreover, $L_{\nabla f}>0$ and $L_{\nabla g}>0$.
{assumption}
The upper-level loss $g$ is convex and bounded below.
Remark 2.1.
The first part of Section 2 implies that the lower-level objective function is strongly convex. The convexity of the upper-level loss $g$ in Section 2 encompasses commonly used loss functions such as mean-squared error (MSE) and multinomial logistic loss [29]. Another popular non-convex loss, peak signal-to-noise ratio (PSNR), is connected to MSE. For other non-convex losses like structural similarity index measure (SSIM) [15], our method and theory remain applicable with slight modifications in the proofs and the line search conditions that we derive.
The gradient of $f$ takes the form below
$\nabla f(\theta)=\nabla(g\circ\hat{x})(\theta)=\partial\hat{x}(\theta)^{T}\nabla g(\hat{x}(\theta)).$ (3)
For all $x\in\mathbb{R}^{n}$, under Section 2 and with the use of the implicit function theorem [3], we can state
$\partial\hat{x}(\theta)^{T}\coloneqq-\nabla_{x\theta}^{2}h(x,\theta)^{T}\nabla_{x}^{2}h(x,\theta)^{-1}.$
Based on Section 2 and Section 2, we have the following useful inequalities, which we will use in later sections. As stated in [30, Lemma 1.2.3], if $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a $L_{\nabla f}$-smooth function, then for any $x,y\in\mathbb{R}^{d}$ we have
$f(y)\leq f(x)+\nabla f(x)^{T}(y-x)+\frac{L_{\nabla f}}{2}\|y-x\|^{2}.$ (4)
Note that (see e.g. [30]) if $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is convex and smooth, then for any $x,y\in\mathbb{R}^{d}$ we have
$f(y)\geq f(x)+\nabla f(x)^{T}(y-x).$ (5)
At each iteration $k\in\mathbb{N}_{0}$, as finding ${\hat{x}(\theta_{k})}$ exactly is not feasible due to the usage of numerical solvers, we solve (2b) inexactly with tolerance $\epsilon_{k}$ to find $\tilde{x}(\theta_{k})$ such that $\|{\tilde{x}(\theta_{k})}-{\hat{x}(\theta_{k})}\|\leq\epsilon_{k}$ using the a posteriori bound $\frac{\|\nabla_{x}h({\tilde{x}(\theta_{k})},\theta_{k})\|}{\mu(\theta_{k})}\leq\epsilon_{k}$. This a posteriori bound is derived from the strong convexity of the lower-level objective function, as explained in [18]. Also, instead of computing $\nabla_{x}^{2}h({\tilde{x}(\theta_{k})},\theta_{k})^{-1}\nabla g(\tilde{x}(\theta_{k}))$, we solve the linear system $\nabla_{x}^{2}h(\tilde{x}(\theta_{k}),\theta_{k})q_{k}=\nabla g(\tilde{x}(\theta_{k}))$ with residual $\delta_{k}$. Then, we denote the inexact gradient (hypergradient) of the upper-level problem (2a) at each iteration $k$ by $z_{k}\coloneqq-\nabla_{x\theta}h({\tilde{x}(\theta_{k})},\theta_{k})^{T}q_{k}$ and the corresponding error by $e_{k}\coloneqq z_{k}-\nabla f(\theta_{k})$.
The following theorem provides a computable a posteriori bound for the error of the approximate hypergradient, which we utilize in the next section.
Theorem 2.2.
([20, Theorem 10]) Suppose Assumption 2 hold. Let us define
$c(x(\theta))\coloneqq\frac{L_{\nabla_{g}}\|\nabla_{x\theta}^{2}h(x(\theta),\theta)\|}{\mu}+L_{H^{-1}}\|\nabla g(x(\theta))\|\|\nabla_{x\theta}^{2}h(x(\theta),\theta)\|+\frac{L_{J}\|\nabla g(x(\theta))\|}{\mu},$
where $L_{H^{-1}}$ is the Lipschitz constant of $\nabla_{x}^{2}h(x,\theta)^{-1}$ in $x$ uniformly for all $\theta$ [^1] From Section 2, it follows that $\nabla_{x}^{2}h^{-1}$ is Lipschitz continuous. Then, at each iteration $k=1,2,\dots$, we have the following a posteriori bound for the inexact gradient:
$\|e_{k}\|=\|z_{k}-\nabla f(\theta_{k})\|\leq c(\tilde{x}(\theta_{k}))\epsilon_{k}+\frac{\|\nabla_{x\theta}^{2}h(\tilde{x}(\theta_{k}),\theta_{k})\|}{\mu}\delta_{k}+\frac{L_{J}L_{\nabla g}}{\mu}\epsilon_{k}^{2}.$ (6)
3 Method of Adaptive Inexact Descent (MAID) and Convergence Analysis
In this section, we propose the Method of Adaptive Inexact Descent (MAID) for solving the problem Section 2, stated in Algorithm 1. This algorithm uses the update
$\theta_{k+1}=\theta_{k}-\alpha_{k}z_{k},\quad k=0,1,\dots,$ (7)
where $z_{k}$ represents the inexact hypergradient at iteration $k$, as computed in Algorithm 2. An appropriate step size $\alpha_{k}>0$ is computed using backtracking. However, conventional backtracking line search rules, such as the Armijo rule [31], require access to an exact function evaluation. This could be available up to machine precision in (2) but would be computationally prohibitively expensive. To address this challenge, we introduce a verifiable line search method that generalises the Armijo rule to the inexact setting utilizes only an approximation of $f$. We control the accuracies to ensure that $-z_{k}$ is a descent direction. Our analysis begins with the Armijo rule in the exact setting, where we have access to $f$ and we can utilize the negative of the exact gradient, that is $-\nabla f(\theta_{k})$, as the descent direction.
The computation of the approximate hypergradient $z_{k}$ in Algorithm 2 employs the implicit function theorem and numerical solvers, following the same procedure as outlined in the previous section, with tolerances $\epsilon_{k}$ and $\delta_{k}$. It checks if the sufficient condition for $-z_{k}$ to be a descent direction for $f$ is satisfied. If not, it shrinks $\epsilon_{k}$ and $\delta_{k}$, recalculates $z_{k}$, and repeats this process until $-z_{k}$ satisfies the requirement. With the approximate hypergradient $z_{k}$ and the tolerance $\epsilon_{k}$ which was used to compute it, Algorithm 1 employs the line search rule $\psi(\alpha_{k})\leq 0$, where $\psi$ is defined in (10), utilizing approximations and bounds for $f(\theta_{k})$ and $\nabla f(\theta_{k})$. If it fails to find a suitable step size and reaches the maximum number of backtracking iterations, it reduces $\epsilon_{k}$, recalculates $z_{k}$, and retries the backtracking process with an increased maximum number of iterations until it successfully finds an appropriate step size and completes the descent update. This backtracking failure can occur due to insufficient iterations to capture a very small step size or $\epsilon_{k}$ that is not small enough.
Algorithm 1 Method of Adaptive Inexact Descent (MAID). Hyperparameters: $\underline{\rho},\overline{\rho}\in(0,1)$ control the reduction and increase of the step size $\alpha_{k}$, respectively; $\underline{\nu},\overline{\nu}\in(0,1)$ govern the reduction and increase of accuracies $\delta_{k}$ and $\epsilon_{k}$; $\max_{\text{BT}}\in\mathbb{N}$ is the maximum number of backtracking iterations.
1:Input $\theta_{0}\in\mathbb{R}^{d}$, accuracies $\epsilon_{0},\delta_{0}>0$, step size $\alpha_{0}>0$.
2:for $k=0,1,\dots$ do
3:     for $j=\max_{\text{BT}},\max_{\text{BT}}+1,\dots$  do
4:         $z_{k},\epsilon_{k},\delta_{k}\leftarrow$ INEXACTGRADIENT($\theta_{k},\epsilon_{k},\delta_{k}$)
5:         for $i=0,1,\dots,j-1$ do
6:              if inexact sufficient decrease $\psi(\alpha_{k})\leq 0$ holds thennormal-$\triangleright$ Lemma 3.7
7:                  Go to line 11$\triangleright$ Backtracking Successful
8:              $\alpha_{k}\leftarrow\underline{\rho}\alpha_{k}$triangleright Adjust the starting step size
9:         $\epsilon_{k}\leftarrow\underline{\nu}\epsilon_{k}$triangleright Backtracking Failed and needs higher accuracy
10:         $\delta_{k}\leftarrow\underline{\nu}\delta_{k}$
11:     $\theta_{k+1}\leftarrow\theta_{k}-\alpha_{k}z_{k}$$\triangleright$ Gradient descent update
12:     $\epsilon_{k+1}\leftarrow\overline{\nu}\epsilon_{k}$triangleright Increasing $\epsilon_{k}$
13:     $\delta_{k+1}\leftarrow\overline{\nu}\delta_{k}$triangleright Increasing $\delta_{k}$
14:     $\alpha_{k+1}\leftarrow\overline{\rho}\alpha_{k}$$\triangleright$ Increasing $\beta_{k}$
Algorithm 2 Calculating an inexact hypergradient, ensuring it is a descent direction. Hyperparameters: $\eta\in(0,1)$ ensures that the computed inexact gradient $z$ is a descent direction; $\underline{\nu}\in(0,1)$ is used to reduce the accuracies $\delta$ and $\epsilon$.
1:Input: $\theta\in\mathbb{R}^{d}$, accuracies $\epsilon,\delta>0$.
2:function InexactGradient($\theta,\epsilon,\delta$)
3:     while True do
4:         Solve lower-level problem to find $\tilde{x}(\theta)$ such that $\|\nabla_{x}h(\tilde{x}(\theta),\theta)\|\leq\epsilon\mu$.
5:         Solve $\nabla_{x}^{2}h(\tilde{x}(\theta),\theta)q=\nabla g(\tilde{x}(\theta))$ with tolerance $\delta$.
6:         Calculate $z=-(\nabla^{2}_{x\theta}h(\tilde{x}(\theta),\theta))^{T}q$.
7:         $\omega\leftarrow$ Calculate upper bound (6) for error $\|e\|$ using Theorem 2.2.
8:         if $\omega\leq(1-\eta)\|z\|$ then
9:              return $z,\epsilon,\delta$
10:         $\delta\leftarrow\underline{\nu}\delta$,
11:         $\epsilon\leftarrow\underline{\nu}\epsilon$.
3.1 Line search with approximate hypergradient and inexact function evaluations
In order to find a suitable step size when we do not have any information about the Lipschitz constant of the gradient, line search methods are usually employed [31]. In Armijo rule, at each iteration $k$, for a given direction $z_{k}$ and a starting step size $\beta_{k}$, the task is to find the smallest $i_{k}\in\mathbb{N}_{0}$ such that
$f(\theta_{k}-\beta_{k}\rho^{i_{k}}z_{k})\leq f(\theta_{k})-\zeta\beta_{k}\rho^{i_{k}}\nabla f(\theta_{k})^{T}z_{k}$ (8)
holds for $10<\rho,\zeta<$ 1. After finding $i_{k}$, we denote step size by $\alpha_{k}\coloneqq\beta_{k}\rho^{i_{k}}$. In our setting, we do not have access to the function $f$ nor its gradient; instead, we only have $g(\tilde{x}(\theta_{k}))$ and an inexact direction $z_{k}$. Therefore, we will introduce a line search condition with inexact gradient and function evaluations to find a suitable step size.
We commence by deriving a condition under which the approximate hypergradient is a descent direction. Subsequently, we establish computable upper and lower bounds for the exact upper-level function by leveraging the inexact components. Moving forward, we introduce our line search.
3.1.1 Descent direction
The following proposition provides a sufficient condition to ensure that $-z_{k}$ is a descent direction for $f$ at $\theta_{k}$, that is $z_{k}^{T}\nabla f(\theta_{k})>0$.
Proposition 3.1.
Let $e_{k}\coloneqq z_{k}-\nabla f(\theta_{k})$ be the error in the approximate hypergradient. If $\|e_{k}\|<\|z_{k}\|$, then $-z_{k}$ is a descent direction for $f$ at $\theta_{k}$.
Proof 3.2.
Since
$z_{k}^{T}\nabla f(\theta_{k})=z_{k}^{T}(z_{k}-e_{k})=\|z_{k}\|^{2}-z_{k}^{T}e_{k},$
$-z_{k}$ is a descent direction if and only if $z_{k}^{T}e_{k}<\|z_{k}\|^{2}.$
Now, from the assumption, by multiplying both sides of the inequality $\|e_{k}\|<\|z_{k}\|$ by $\|z_{k}\|>0$ we have
$\|e_{k}\|\|z_{k}\|<\|z_{k}\|^{2}.$
Moreover, using Cauchy–Schwarz, we have
$z_{k}^{T}e_{k}\leq\|z_{k}\|\|e_{k}\|<\|z_{k}\|^{2},$
thus $-z_{k}$ is a descent direction.
Note that the condition ${\|e_{k}\|}<\|z_{k}\|$ in Proposition 3.1 must hold uniformly to demonstrate the convergence of a descent method [31, Theorem 3.2]. Therefore, instead, we will assume the stronger condition ${\|e_{k}\|}\leq(1-\eta){\|z_{k}\|}$ for $10<\eta<$ 1.
In addition, due to the unavailability of $\nabla f(\theta_{k})$, we establish a computable bound for $z_{k}^{T}\nabla f(\theta_{k})$ to facilitate our analysis.
Lemma 3.3.
Let ${\|e_{k}\|}\leq(1-\eta){\|z_{k}\|}$. Then, $\eta\|z_{k}\|^{2}\leq\nabla f(\theta_{k})^{T}z_{k}.$
Proof 3.4.
Combining Cauchy–Schwarz and ${\|e_{k}\|}\leq(1-\eta){\|z_{k}\|}$ yields $e_{k}^{T}z_{k}\leq\|z_{k}\|^{2}-\eta\|z_{k}\|^{2}.$ Since $\nabla f(\theta_{k})^{T}z_{k}=\|z_{k}\|^{2}-e_{k}^{T}z_{k}$, we find the desired bound.
We now derive computable estimates for $f(\theta_{k})$ which will be used to establish our alternative sufficient decrease condition.
3.1.2 Verifiable line search with approximate hypergradient
The following lemma derives computable lower and upper bounds for $f$.
Lemma 3.5.
Let $g$ be $L_{\nabla g}$-smooth and convex. We have the following lower and upper bounds for $f(\theta_{k})=g({\hat{x}(\theta_{k})})$,
$g({\hat{x}(\theta_{k})})\leq g({\tilde{x}(\theta_{k})})+\|\nabla g({\tilde{x}(\theta_{k})})\|\epsilon_{k}+\frac{L_{\nabla_{g}}}{2}\epsilon_{k}^{2},$ (9a)
$g({\hat{x}(\theta_{k})})\geq g({\tilde{x}(\theta_{k})})-\|\nabla g({\tilde{x}(\theta_{k})})\|\epsilon_{k}.$ (9b)
Proof 3.6.
Since $g$ is $L_{\nabla g}$-smooth, using the inequality (4) we have
$g({\tilde{x}(\theta_{k})})-g({\hat{x}(\theta_{k})})-\nabla g(\tilde{x}(\theta_{k}))^{T}(\tilde{x}(\theta_{k})-\hat{x}(\theta_{k}))\leq\frac{L_{\nabla g}}{2}\|\tilde{x}(\theta_{k})-\hat{x}(\theta_{k})\|^{2}.$
Now, by rearranging and utilizing the Cauchy-Schwarz inequality, we derive
$g({\tilde{x}(\theta_{k})})-g({\hat{x}(\theta_{k})})\leq\frac{L_{\nabla g}}{2}\|\tilde{x}(\theta_{k})-\hat{x}(\theta_{k})\|^{2}+\|\nabla g(\tilde{x}(\theta_{k}))\|\|(\tilde{x}(\theta_{k})-\hat{x}(\theta_{k}))\|.$
Since $\|\tilde{x}(\theta_{k})-\hat{x}(\theta_{k})\|\leq\epsilon_{k}$, the inequality (9a) holds.
Furthermore, since $g$ is convex, using the inequality (5), we can write
$f(\theta_{k})=g({\hat{x}(\theta_{k})})\geq g({\tilde{x}(\theta_{k})})+\nabla g({\tilde{x}(\theta_{k})})^{T}({\hat{x}(\theta_{k})}-{\tilde{x}(\theta_{k})}).$
Following the same argument as before we derive the desired inequality (9b).
The following lemma employs the bounds (9a) and (9b) to provide a line search rule such that all of its components are accessible.
Lemma 3.7.
Let $\lambda\in\mathbb{R}$ and suppose that $g$ is $L_{\nabla g}$-smooth and convex, and denote
$\psi(\alpha_{k})\coloneqq g(\tilde{x}(\theta_{k+1}))+\|\nabla g(\tilde{x}(\theta_{k+1}))\|\epsilon_{k+1}+\frac{L_{\nabla_{g}}}{2}\epsilon_{k+1}^{2}-g(\tilde{x}(\theta_{k}))+\|\nabla g(\tilde{x}(\theta_{k}))\|\epsilon_{k}+\lambda\alpha_{k}\|z_{k}\|^{2}.$ (10)
If the backtracking line search condition $\psi(\alpha_{k})\leq 0$ is satisfied, then the sufficient descent condition $f(\theta_{k+1})-f(\theta_{k})\leq-\lambda\alpha_{k}\|z_{k}\|^{2}$ holds.
Proof 3.8.
Using (9a) for $\theta_{k+1}$ and (9b) for $\theta_{k}$, we have
$f(\theta_{k+1})-f(\theta_{k})\leq g(\tilde{x}(\theta_{k+1}))+\|\nabla g(\tilde{x}(\theta_{k+1}))\|\epsilon_{k+1}+\frac{L_{\nabla_{g}}}{2}\epsilon_{k+1}^{2}-g(\tilde{x}(\theta_{k}))+\|\nabla g(\tilde{x}(\theta_{k}))\|\epsilon_{k}.$
The inequality above together with $\psi(\alpha_{k})\leq 0$ imply $f(\theta_{k+1})-f(\theta_{k})\leq-\lambda\alpha_{k}\|z_{k}\|^{2}$ as required.
Lemma 3.7 provides a backtracking line search condition for sufficient decrease. We adopt the inequality $\psi(\alpha_{k})\leq 0$ as our inexact backtracking line search rule in Algorithm 1, where, for $\alpha_{k}=\underline{\rho}^{i_{k}}\beta_{k}$, $i_{k}\geq 0$ is the smallest integer satisfying the inequality. Note that this is a practical line search as we can compute all of its components.
Now, since the components of the line search rule $\psi(\alpha_{k})\leq 0$ depend on $\epsilon_{k}$, we investigate the accuracy $\epsilon_{k}$ for which a step size $\alpha_{k}$ exists that satisfies $\psi(\alpha_{k})\leq 0$ a priori.
3.1.3 Existence of a step size in the line search
Lemma 3.9.
Let $\lambda\in\mathbb{R}$, $\eta\leq 1$, and ${\|e_{k}\|}\leq(1-\eta){\|z_{k}\|}$. Denote
$w_{k}\coloneqq\|\nabla g(\tilde{x}(\theta_{k}))\|+\|\nabla g(\tilde{x}(\theta_{k+1}))\|,$ (11)
$\bar{\epsilon}_{k}\coloneqq\max\{\epsilon_{k},\epsilon_{k+1}\}$, and $\phi(\alpha_{k})\coloneqq 2w_{k}\bar{\epsilon}_{k}+{L_{\nabla_{g}}}\bar{\epsilon}_{k}^{2}+(\alpha_{k}^{2}\frac{L_{\nabla f}}{2}+(\lambda-\eta)\alpha_{k})\|z_{k}\|^{2}$. Then,
$\psi(\alpha_{k})\leq\phi(\alpha_{k}).$ (12)
Proof 3.10.
Since $f(\theta)=g(\hat{x}(\theta))$ is $L_{\nabla f}-$smooth, applying (4) to $g(\hat{x}(\theta_{k+1}))=g(\hat{x}(\theta_{k}-\alpha_{k}z_{k}))$ in (9b) for $\theta_{k+1}$ gives
$g(\tilde{x}(\theta_{k+1}))$ $\leq g(\hat{x}(\theta_{k+1}))+\|\nabla g(\tilde{x}(\theta_{k+1}))\|\epsilon_{k+1}$
$\leq g(\hat{x}(\theta_{k}))-\alpha_{k}\nabla f(\theta_{k})^{T}z_{k}+\alpha_{k}^{2}\frac{L_{\nabla f}}{2}\|z_{k}\|^{2}+\|\nabla g(\tilde{x}(\theta_{k+1}))\|\epsilon_{k+1}.$
Leveraging the bounds (9a) for $g(\hat{x}(\theta_{k}))$ and Lemma 3.3 for $\nabla f(\theta_{k})^{T}z_{k}$, we derive
$g(\tilde{x}(\theta_{k+1}))$ $\leq g(\tilde{x}(\theta_{k}))+\|\nabla g(\tilde{x}(\theta_{k}))\|\epsilon_{k}+\frac{L_{\nabla g}}{2}\epsilon_{k}^{2}+\left(\alpha_{k}^{2}\frac{L_{\nabla f}}{2}-\alpha_{k}\eta\right)\|z_{k}\|^{2}+\|\nabla g(\tilde{x}(\theta_{k+1}))\|\epsilon_{k+1}.$
Adding $\|\nabla g(\tilde{x}(\theta_{k+1}))\|\epsilon_{k+1}+\frac{L_{\nabla_{g}}}{2}\epsilon_{k+1}^{2}-g(\tilde{x}(\theta_{k}))+\|\nabla g(\tilde{x}(\theta_{k}))\|\epsilon_{k}+\lambda\alpha_{k}\|z_{k}\|^{2}$ to both sides we obtain
$\psi(\alpha_{k})\leq 2\|\nabla g(\tilde{x}(\theta_{k}))\|\epsilon_{k}+2\|\nabla g(\tilde{x}(\theta_{k+1}))\|\epsilon_{k+1}+\frac{L_{\nabla g}}{2}\left(\epsilon_{k}^{2}+\epsilon_{k+1}^{2}\right)+\left(\alpha_{k}^{2}\frac{L_{\nabla f}}{2}+\alpha_{k}(\lambda-\eta)\right)\|z_{k}\|^{2}.$
Now, utilizing $\bar{\epsilon}_{k}=\max\{\epsilon_{k},\epsilon_{k+1}\}$, and $w_{k}=\|\nabla g(\tilde{x}(\theta_{k}))\|+\|\nabla g(\tilde{x}(\theta_{k+1}))\|$, we derive the desired result.
Lemma 3.11.
Assume $\|z_{k}\|\neq 0$, $\lambda<\eta$, and ${\|e_{k}\|}\leq(1-\eta){\|z_{k}\|}$. Considering $w_{k}$ and $\bar{\epsilon}$ as defined in Lemma 3.9 and denoting
$s_{k}\coloneqq\frac{1}{L_{\nabla g}}{\left(\sqrt{w_{k}^{2}+\frac{L_{\nabla g}}{2L_{\nabla f}}(\eta-\lambda)^{2}\|z_{k}\|^{2}}-w_{k}\right)},$ (13)
for all $\bar{\epsilon}_{k}\in[0,s_{k}),$ there exist $0\leq\underline{\alpha}_{k}<\overline{\alpha}_{k}$ such that for all $\alpha_{k}\in[\underline{\alpha}_{k},\overline{\alpha}_{k}]$ the sufficient decrease condition $\psi(\alpha_{k})\leq 0$ holds. Moreover, denoting
$\hat{s}_{k}\coloneqq\sqrt{(\eta-\lambda)^{2}-\frac{2L_{\nabla f}(2w_{k}\bar{\epsilon}_{k}+L_{\nabla g}\bar{\epsilon}_{k}^{2})}{\|z_{k}\|^{2}}}>0,$ (14)
the interval may be defined by
$\underline{\alpha}_{k}=\frac{1}{L_{\nabla f}}\left({\eta-\lambda}-\hat{s}_{k}\right),\quad\overline{\alpha}_{k}=\frac{1}{L_{\nabla f}}\left({\eta-\lambda}+\hat{s}_{k}\right).$ (15)
Note that $\lim_{\bar{\epsilon_{k}}\to 0}\underline{\alpha}_{k}=0$, $\lim_{\bar{\epsilon_{k}}\to 0}\overline{\alpha}_{k}=\frac{2(\eta-\lambda)}{L_{\nabla f}}>0$, and $\hat{s}_{k}$ is monotonically decreasing in $\bar{\epsilon}_{k}$.
Proof 3.12.
From Lemma 3.9 it follows that $\phi(\alpha_{k})\leq 0$ implies $\psi(\alpha_{k})\leq 0$. Note that $\phi(\alpha_{k})$ is quadratic in $\alpha_{k}$ with positive leading coefficient. The roots of $\phi$ are given by $\underline{\alpha}_{k}$ and $\overline{\alpha}_{k}$ (15). These roots are real if
$(\eta-\lambda)^{2}\|z_{k}\|^{2}-2L_{\nabla f}(2w_{k}\bar{\epsilon}_{k}+L_{\nabla g}\bar{\epsilon}_{k}^{2})\geq 0.$ (16)
Following a similar argument, the inequality (16) holds if $\bar{\epsilon}_{k}\leq s_{k}$, given that the left-hand side of (16) is quadratic in $\bar{\epsilon}_{k}$ with a negative leading coefficient and roots $s_{k}$ and $-s_{k}$. Moreover, $s_{k}>0$ by definition, as $\|z_{k}\|\neq 0$. Since $\bar{\epsilon}_{k}\geq 0$, we have $\bar{\epsilon}_{k}\in[0,s_{k})$. Referring to (16) and the definition of $\hat{s}_{k}$, we conclude that $0<\hat{s}_{k}\leq{\eta-\lambda}$, implying $0\leq\underline{\alpha}_{k}<\overline{\alpha}_{k}$.
As $s_{k}$ (13) depends on ${\epsilon}_{k}$ and ${\epsilon}_{k+1}$, and thus on $\bar{\epsilon}_{k}$, we need the following lemmas to show that for $\bar{\epsilon}_{k}$ sufficiently small, the inclusion $\bar{\epsilon}_{k}\in[0,s_{k})$ holds a priori for any $z_{k}$, $\tilde{x}(\theta_{k})$, and $\tilde{x}(\theta_{k+1})$ satisfying the conditions of computing the inexact gradient and the solution of the lower-level problem. It shows that if $\theta_{k}$ is non-stationary, there exists a nonempty positive range of accuracies for which we can find a nonempty positive range of step sizes satisfying the line search condition $\psi(\alpha_{k})\leq 0$. For simplicity, we set $\delta_{k}=\bar{\epsilon}_{k}$ for the remainder of this section, however, the general case holds with a similar argument.
Lemma 3.13.
Let $\|\nabla f(\theta_{k})\|>0$, $s_{k}$ be as defined in (13), and $\lambda\neq\eta$, then $\lim_{\bar{\epsilon_{k}}\to 0}s_{k}>0$.
Proof 3.14.
As $\lim_{\bar{\epsilon_{k}}\to 0}\tilde{x}(\theta_{k})=\hat{x}(\theta_{k})$, $\lim_{\bar{\epsilon_{k}}\to 0}\tilde{x}(\theta_{k+1})=\hat{x}(\theta_{k+1})$, and $\nabla g$ is continuous, we have $\lim_{\bar{\epsilon_{k}}\to 0}w_{k}=\bar{w}_{k},$ for some $\bar{w}_{k}\geq 0$. Moreover, since $\lim_{\bar{\epsilon_{k}}\to 0}z_{k}=\nabla f(\theta_{k})$, $\|\nabla f(\theta_{k})\|>0$, $\lambda\neq\eta$, and $L_{\nabla f}>0$,
$\lim_{\bar{\epsilon_{k}}\rightarrow 0}s_{k}=\frac{1}{L_{\nabla g}}{\left(\sqrt{\bar{w}_{k}^{2}+\frac{L_{\nabla g}}{2L_{\nabla f}}(\eta-\lambda)^{2}\|\nabla f(\theta_{k})\|^{2}}-\bar{w}_{k}\right)}>0.$
as desired.
Now, we need the following lemma to show that for $\bar{\epsilon}_{k}$ small enough, the inclusion $\bar{\epsilon}_{k}\in[0,s_{k})$ holds a priori.
Lemma 3.15.
Let $\{a_{j}\}_{j=0}^{\infty}\subseteq[0,\infty)$ and $\{b_{j}\}_{j=0}^{\infty}\subseteq[0,\infty)$ be sequences, with $a_{j}\rightarrow a>0$ and $b_{j}\rightarrow 0$, respectively. Then, there exists $J\in\mathbb{N}$ such that for all $j\geq J$, $b_{j}\in[0,a_{j}]$.
Proof 3.16.
[…]
Corollary 3.17.
Let $\|\nabla f(\theta_{k})\|>0$, $\bar{\epsilon}_{k}\to 0$, and $10<\lambda<\eta<$ 1. Then, there exists $\epsilon>0$ such that if $\bar{\epsilon}_{k}\leq\epsilon$, then $\bar{\epsilon}_{k}<s_{k}$.
That means, if the accuracy is small enough, there exist a range of step sizes satisfying $\psi(\alpha_{k})\leq 0$. Note that $\epsilon$ is independent of $z_{k}$, $\tilde{x}(\theta_{k})$, and $\hat{x}(\theta_{k})$.
Proof 3.18.
The assertion follows by setting $b_{j}=(\bar{\epsilon}_{k})_{j}$ and $a_{j}=(s_{k})_{j}$ using Lemma 3.13 and Lemma 3.15.
Lemma 3.19.
Let $\lambda<\eta$, $10<\underline{\rho}<$ 1, and $\underline{\alpha}_{k}$ and let $\overline{\alpha}_{k}$ be as defined in (15).
  • If $\bar{\epsilon}_{k}\geq 0$ be sufficiently small, then there exists $j_{k}\in\mathbb{N}_{0}$ such that $\underline{\alpha}_{k}\leq\beta_{k}\underline{\rho}^{j_{k}}\leq\overline{\alpha}_{k}$, i.e. $\psi(\alpha_{k})\leq 0$ holds for $\alpha_{k}=\beta_{k}\underline{\rho}^{j_{k}}$.
  • If $i_{k}>0$ is the smallest integer for which $\psi(\alpha_{k})\leq 0$ holds, then $\alpha_{k}=\beta_{k}\underline{\rho}^{i_{k}}>\underline{\rho}\overline{\alpha}_{k}$.
Proof 3.20.
By Lemma 3.11, for all $\bar{\epsilon}_{k}\in[0,s_{k})$, $\overline{\alpha}_{k}>0$. Let $\bar{\epsilon}_{k}\in[0,s_{k})$; since $\lim_{j\to\infty}\beta_{k}\underline{\rho}^{j}=0$, there exists an $j_{k}\in\mathbb{N}_{0}$ such that for all $j\geq j_{k}$, $\beta_{k}\underline{\rho}^{j}\leq\overline{\alpha}_{k}$. Referring again to Lemma 3.11, we know that $\lim_{\bar{\epsilon}_{k}\to 0}\underline{\alpha}_{k}=0$. Therefore, there exists $\epsilon\in[0,s_{k})$ such that for all $\bar{\epsilon}_{k}\leq\epsilon$, $\underline{\alpha}_{k}\leq\beta_{k}\underline{\rho}^{j_{k}}$. Additionally, $\overline{\alpha}_{k}$ as a function of $\bar{\epsilon}_{k}$ is decreasing on the domain $[0,s_{k})$. Hence, $\beta_{k}\underline{\rho}^{j_{k}}\leq\overline{\alpha}_{k}$ holds for $\bar{\epsilon}_{k}\leq\epsilon$.
As $i_{k}>0$ be the smallest integer for which $\psi(\alpha_{k})\leq 0$ holds, then $\alpha_{k}=\beta_{k}\underline{\rho}^{i_{k}}\leq\overline{\alpha}_{k}$. If $\alpha_{k}\leq\underline{\rho}\overline{\alpha}_{k}$, then $\beta_{k}\underline{\rho}^{i_{k}-1}\leq\overline{\alpha}_{k}$, which contradicts with the assumption that $i_{k}$ is the smallest integer for which $\psi(\alpha_{k})\leq 0$ holds. Hence, $\beta_{k}\underline{\rho}^{i_{k}}\leq\underline{\rho}\overline{\alpha}_{k}$
Based on Lemma 3.19, we can guarantee the existence of a suitable step size during the backtracking line search process if $\bar{\epsilon}_{k}$ is small enough.
3.2 Convergence Analysis
Prior to discussing the convergence theorem, we shall introduce the following auxiliary lemma.
Lemma 3.21.
Let $\epsilon>0$ and $0<\underline{c}<1<\overline{c}$. Further, let $\{a_{k}\}_{k=0}^{\infty}\subseteq[0,\infty)$ be the sequence
$a_{k+1}=\begin{cases}\overline{c}a_{k},&a_{k}<\epsilon\\ \underline{c}^{i_{k}}a_{k},&a_{k}\geq\epsilon\end{cases},$
and $i_{k}$ be the smallest non-negative integer such that $\underline{c}^{i_{k}}a_{k}<\epsilon$. Then, for all $k\geq 0$, $a_{k}\geq\min\{\underline{c}\epsilon,a_{0}\}$.
Proof 3.22.
Assume, by induction, that for all $k=0,\dots,t$, $a_{t}\geq\min\{\underline{c}\epsilon,a_{0}\}$. Now, consider $a_{t+1}$. If $a_{t}<\epsilon$, by definition, $a_{t+1}=\overline{c}a_{t}\geq\overline{c}\min\{\underline{c}\epsilon,a_{0}\}>\min\{\underline{c}\epsilon,a_{0}\}$. Otherwise, $a_{t}\geq\epsilon$. If $\underline{c}\epsilon\geq\underline{c}^{i_{t}}a_{t}$, then $\epsilon\geq\underline{c}^{i_{t}-1}a_{t}$, which contradicts the definition of $i_{t}$. Hence, $\underline{c}\epsilon<\underline{c}^{i_{t}}a_{t}$, which implies $a_{t+1}=\underline{c}^{i_{t}}a_{t}>\underline{c}\epsilon\geq\min\{\underline{c}\epsilon,a_{0}\}$.
Lemma 3.23.
Let Assumptions 2 and 2 hold, $\beta_{0}>0$, and $0<\underline{\rho}<1<\overline{\rho}$. Let $i_{k}$ be the smallest non-negative integer such that $\alpha_{k}=\underline{\rho}^{i_{k}}\beta_{k}\in[\underline{\alpha}_{k},\overline{\alpha}_{k}]$ and $\alpha_{k}>\underline{\rho}\overline{\alpha}_{k}$, if $i_{k}>0$. Further, let $\beta_{k+1}=\overline{\rho}\alpha_{k}$, $\|z_{k}\|\neq 0$, and ${\|e_{k}\|}\leq(1-\eta){\|z_{k}\|}$. Denoting
$\tau\coloneqq\min\left\{\underline{\rho}\left(\frac{\eta-\lambda}{L_{\nabla f}}\right),\underline{\rho}\left(\frac{2(\eta-\lambda)}{L_{\nabla f}}\right),\beta_{0}\ \right\},$ (17)
we have $\alpha_{k}\geq\tau$.
Proof 3.24.
If $i_{k}>0$, since $\alpha_{k}=\underline{\rho}^{i_{k}}\beta_{k}$, using Lemma 3.19 and (15) we have
$\alpha_{k}>\underline{\rho}\overline{\alpha}_{k}\geq\underline{\rho}\left(\frac{\eta-\lambda}{L_{\nabla f}}\right).$ (18)
Moreover, from Lemma 3.11, (15) yields
$\alpha_{k}\leq\overline{\alpha}_{k}\leq\frac{2(\eta-\lambda)}{L_{\nabla f}}.$ (19)
If $i_{k}=0$, then $\alpha_{k}=\beta_{k}$. Considering the update $\beta_{k+1}=\overline{\rho}\alpha_{k}$, utilizing Lemma 3.21 by setting $a_{k}=\beta_{k}$, $\underline{c}=\underline{\rho}$, $\overline{c}=\overline{\rho}$, and $\epsilon=\frac{2(\eta-\lambda)}{L_{\nabla f}}$, based on (19), we derive $\beta_{k}\geq\min\{\underline{\rho}\left(\frac{2(\eta-\lambda)}{L_{\nabla f}}\right),\beta_{0}\}$. Taking $\tau$ as defined in (17) yields the desirable result.
Theorem 3.25.
Let the assumptions of Lemma 3.23 hold and $10<\lambda<\eta<$ 1. We have
$\lim_{k\rightarrow\infty}\|z_{k}\|=0.$
Proof 3.26.
Lemma 3.7 and the backtracking line search rule $\psi(\alpha_{k})\leq 0$ yield
$f(\theta_{k+1})-f(\theta_{k})\leq-\lambda\alpha_{k}\|z_{k}\|^{2}.$ (20)
Utilizing Lemma 3.23, since $10<\lambda<\eta<$ 1, we have $\tau>0$, and recursively from the inequality (20) we can conclude
$\begin{split}\sum_{k=1}^{K}\lambda\tau\|z_{k}\|^{2}\leq f(\theta_{1})-f(\theta_{K+1}).\end{split}$ (21)
As $f$ is bounded below and $\lambda\tau>0$ we conclude $\lim_{k\rightarrow\infty}\|z_{k}\|=0$.
Corollary 3.27.
Under the assumptions of Theorem 3.25 we have
$\lim_{k\rightarrow\infty}\|\nabla f(\theta_{k})\|=0.$
Proof 3.28.
Since ${\|e_{k}\|}\leq(1-\eta)\|z_{k}\|$, utilizing Theorem 3.25 we have $\lim_{k\rightarrow\infty}\|z_{k}\|=0$, which implies $\lim_{k\rightarrow\infty}\|e_{k}\|=0$. Hence
$\lim_{k\rightarrow\infty}\|\nabla f(\theta_{k})\|\leq\lim_{k\rightarrow\infty}(\|e_{k}\|+\|z_{k}\|)=0,$
as required.
Note that, at each iteration $k\in\mathbb{N}_{0}$, based on Proposition 3.1 and Theorem 3.25, the approximate hypergradient should satisfy the inequality ${\|e_{k}\|}\leq(1-\eta){\|z_{k}\|}$. Otherwise, we can decrease the error in the hypergradient $\|e_{k}\|$ by means of multiplying $\epsilon_{k}$ and $\delta_{k}$ by a factor $10<\underline{\rho}<$ 1 as we do in the steps 10 and 11, then restarting the calculation of $z_{k}$ using Algorithm 2 until ${\|e_{k}\|}\leq(1-\eta){\|z_{k}\|}$ holds. Moreover, the steps of Algorithm 1 are designed in a way to satisfy both the required accuracy for the line search $\psi(\alpha_{k})\leq 0$. MAID also allows $\epsilon_{k}$ and $\delta_{k}$ to increase in steps 12 and 13 if they do not decrease in the subsequent iteration and remain suitable for the line search condition, which matches with $\bar{\epsilon}_{k}$ as mentioned in Lemma 3.9 and Lemma 3.11. This process in Algorithm 1 and Algorithm 2 provides us with an adaptive way of choosing the accuracy.
Proposition 3.29.
If $\|\nabla f(\theta_{k})\|>0$, the loop at line 3 in Algorithm 2 terminates in finite time by finding a descent direction $-z_{k}$.
Proof 3.30.
From the bound (6) for the upper bound $\omega$ of the error $\|e_{k}\|=\|z_{k}-\nabla f(\theta_{k})\|$, we have $\lim_{\epsilon_{k},\delta_{k}\to 0}\omega=0$. On the other hand, $\lim_{\epsilon_{k},\delta_{k}\to 0}\|z_{k}\|=\|\nabla f(\theta_{k})\|>0$. Since $\epsilon_{k}\to 0$ and $\delta_{k}\to 0$ according to lines 11 and 10 of Algorithm 2, respectively, and utilizing Lemma 3.15, we conclude that the inequality $\omega\leq(1-\eta)\|z_{k}\|$ holds for $\epsilon_{k}$ and $\delta_{k}$ sufficiently small.
Proposition 3.31.
The loop at line 3 in Algorithm 1 terminates in finite time by successfully performing the update $\theta_{k+1}=\theta_{k}-\underline{\rho}^{i_{k}}\beta_{k}z_{k}$ for some $i_{k}\in\mathbb{N}_{0}$.
Proof 3.32.
If the inner loop at the line 5 fails to find a step size within the given maximum number of iterations $j$, two scenarios may occur. Firstly, if $\epsilon_{k}$ and $\delta_{k}$ are not sufficiently small, they are decreased in steps 9 and 10, followed by updating the descent direction $z_{k}$ and a possible further decrease in accuracies $\epsilon_{k}$ and $\delta_{k}$ in step 4 as a part of the next iteration of the outer loop at line 3. This sequence of accuracy values will eventually become small enough to satisfy the assumptions in Corollary 3.17 and Lemma 3.19. Consequently, the required result holds.
Secondly, if $\underline{\rho}^{j-1}\beta_{k}>\overline{\alpha}_{k}$ and the backtracking process is unsuccessful, then the maximum number of backtracking iterations $j$ goes to infinity during the iterations of the outer loop 3, and $\underline{\rho}^{j}\beta_{k}\to 0$. Since $\overline{\alpha}_{k}>0$, this implies that there exists $j_{k}\in\mathbb{N}_{0}$ such that for all $i_{k}\geq j_{k}$, $\underline{\rho}^{i_{k}}\beta_{k}\leq\overline{\alpha}_{k}$. Therefore, the loop 3 terminates in a finite process.
Theorem 3.33.
Under Assumptions 2 and 2, the iterates $\theta_{k}$ of Algorithm 1 satisfy
$\lim_{k\rightarrow\infty}\|\nabla f(\theta_{k})\|=0.$
Proof 3.34.
From Proposition 3.31 we know that the loop at line 3 terminates after a finite number of iterations and finds a suitable step size $\alpha_{k}$ satisfying the line search condition $\psi(\alpha_{k})\leq 0$. Moreover, the update of the initial step size in line 14 fulfills $\beta_{k+1}=\overline{\rho}\alpha_{k}$. Furthermore, $z_{k}$ computed in Algorithm 2 satisfies ${\|e_{k}\|}\leq(1-\eta){\|z_{k}\|}$. Hence, all the assumptions of Theorem 3.25 are satisfied by Algorithm 1 and convergence follows from Corollary 3.27.
4 Numerical Experiments
In this section, we present the numerical results of our proposed algorithm MAID (Algorithm 1) and compare it with HOAG [33] and the dynamic derivative-free algorithm DFO-LS [18]. For HOAG, we take the adaptive step size strategy used in the numerical experiments of [33] with accuracies either following a geometric sequence $\epsilon_{k}=0.9^{k}\epsilon_{0}$, a quadratic sequence $\epsilon_{k}={k^{-2}}\epsilon_{0}$, or a cubic sequence $\epsilon_{k}={k^{-3}}\epsilon_{0}$ for $k=1,2,\dots$. The implementation of HOAG [^2] https://github.com/fabianp/hoag and the Fast Iterative Shrinkage-Thresholding Algorithm (FISTA) variant of DFO-LS [^3] https://github.com/lindonroberts/inexact_dfo_bilevel_learning is based on the available code from the corresponding papers [33] and [18], respectively. We commit to making the codes of our implementations available on GitHub upon the acceptance of the paper. Moreover, we use warm-starting in the lower-level iterations, which means we save and use the lower-level solution as the initial point for the next call of the lower-level solver. This strategy is also utilized in DFO-LS and HOAG. For the constant $L_{H^{-1}}$ in the bound (6) used in Algorithm 2, we approximate it as
$\frac{\|\nabla_{x}^{2}h(\tilde{x}(\theta),\theta)^{-1}\nabla g(\tilde{x}(\theta))-\nabla_{x}^{2}h(\tilde{x}(\theta),\theta)^{-1}u\|}{\|\nabla g(\tilde{x}(\theta))-u\|},$
where $u\sim\mathcal{N}(0,1)$ is a random perturbation of $\nabla g(\tilde{x}(\theta))$. We then store the maximum value of this approximation across the upper-level iterations. In this approach, since we have already calculated $\nabla_{x}^{2}h(\tilde{x}(\theta),\theta)^{-1}\nabla g(\tilde{x}(\theta))$ to compute the hypergradient, we possess an efficient approximation that improves by taking its maximum across upper-level iterations. We use a similar process for estimating $L_{J}$. We include these calculations in our lower-level computational cost.
4.1 Validating the performance with known optimal parameters
Similar to [20], we first consider a simple linear least-squares problem taken from [25]:
$\min_{\theta\in\mathbb{R}^{10}}f(\theta)$ $\coloneqq\|A_{1}\hat{x}(\theta)-b_{1}\|^{2},$ (22a)
$s.t.\quad\hat{x}(\theta)$ $\coloneqq\arg\min_{x\in\mathbb{R}^{10}}\|A_{2}x+A_{3}\theta-b_{2}\|^{2},$ (22b)
where $A_{i}\in\mathbb{R}^{1000\times 10}$ have random entries uniformly distributed between 0 and 1, and $b_{i}\in\mathbb{R}^{1000}$ are given by $b_{1}=A_{1}\hat{x}_{1}+0.01y_{1}$ and $b_{2}=A_{2}\hat{x}_{2}+A_{3}\bar{\theta}+0.01y_{2}$, where $\hat{x}_{1},\hat{x}_{2},$ and $\bar{\theta}\in\mathbb{R}^{10}$ have i.i.d. entries uniformly distributed between 0 and 1, and $y_{1},y_{2}\in\mathbb{R}^{1000}$ are independent standard Gaussian vectors with entries from $\mathcal{N}(0,1)$. For our experiments, we pick $\theta_{0}$ as the vector of all ones. In this problem, we have the capability to analytically compute $\hat{x}(\theta)$, $\nabla f(\theta)$, and all of the Lipschitz constants ($L_{\nabla f}$, $L_{\nabla g}$, $L_{H^{-1}}$, and $L_{J}$), along with determining the optimal $\theta^{*}$. Therefore, this problem serves as a benchmark to evaluate the performance of our inexact algorithm.
For the experimental evaluation, we have chosen several configurations for the parameters in our algorithm. Specifically, these encompass $\epsilon_{0}=\delta_{0}=10^{-1}$, $\epsilon_{0}=\delta_{0}=10^{-3}$, $\epsilon_{0}=\delta_{0}=10^{-5}$. Moreover, in MAID, we set the hyperparameters $\overline{\rho}=\frac{10}{9}$, $\underline{\rho}=0.5$, $\overline{\nu}=1.25$, and $\underline{\nu}=0.5$. Across all instances of fixed accuracy, the line search mechanism of MAID ($\psi(\alpha_{k})\leq 0$) is employed to determine the step size. Subsequently, we employ FISTA adapted to strongly convex problems to solve the lower-level problem [13].
Since the main computational cost involved in solving bilevel optimization problems lies in solving the lower-level problems, we have imposed a cap on the total number of lower-level iterations. Furthermore, recognizing the computational analogy between the cost of each CG iteration in the IFT approach, dominated by a single Hessian-vector product, and each iteration of the lower-level solver (e.g., FISTA), dominated by a single gradient evaluation, we treat CG iterations with comparable significance to that of an iteration of the lower-level solver. Consequently, the amalgamation of lower-level solver iterations and CG iterations is regarded as constituting the cumulative computational cost of the lower-level phase. For solving the problem (4.1), we set a termination budget of $1.5\times 10^{5}$ lower-level iterations.
Figure 1: The value of $\epsilon$ per upper-level iteration, alongside the performance and computational cost of MAID, is illustrated across various settings for solving the problem (4.1). All MAID configurations achieve lower loss values compared to the high fixed accuracy, and they do so at a lower computational cost. Furthermore, they consistently converge to the same range of the required accuracy $\epsilon=\delta$, regardless of the initial choices of $\epsilon_{0}$.
Figure 0(a) illustrates the determination of the necessary values for $\epsilon$ in solving the problem (4.1) through the MAID algorithm. As shown in Fig. 0(a), irrespective of the initial $\epsilon_{0}=\delta_{0}$ as the accuracy of the lower-level solver and the linear solver, respectively, MAID adaptively reduces $\epsilon$ and $\delta$, ultimately opting for accuracies around $2\times 10^{-5}$.
Fig. 0(b) illustrates the computational advantages of adaptivity in MAID compared to fixed low and high accuracies, with the main computational costs stemming from lower-level and linear solver iterations. Setting the low fixed accuracy to $\epsilon=10^{-1}$ and $\epsilon=10^{-3}$ results in loss stalling after a few upper-level iterations due to insufficient accuracy for line search progression. Conversely, using the higher fixed accuracy of $\epsilon=10^{-5}$ yields a higher loss than dynamic configurations, due to exhausting the total lower-level iteration budget. This underscores the importance of the adaptive approach in MAID, as the suitable fixed accuracy is not known a priori; it may be too small, leading to failure (e.g., $\epsilon=10^{-1}$), or require significantly higher computational costs for progress (e.g., $\epsilon=10^{-5}$). Overall, dynamic configurations successfully solve problem (4.1) with lower computational costs compared to a fixed accuracy of $10^{-5}$, regardless of the starting accuracy.
Figure 2: The range of error in upper-level function evaluation given error bounds (9b) and (9a), inexact loss given the approximated lower-level solution, and exact loss evaluation $f(\theta)=g(\hat{x}(\theta))$.
While in Fig. 0(b) we plot the optimality gap, it is also possible to compute the exact solution of the lower-level problem in the specific problem (4.1). This enables the calculation of the precise loss at each upper-level iteration. This ability allows us to verify both the sufficient decrease condition and the progress of MAID. To facilitate this verification, we present the exact loss alongside the loss obtained using MAID with $\epsilon_{0}=\delta_{0}=1$ in Fig. 2. Additionally, this figure visually represents the bounds for inexact loss evaluation, as defined by (9b) and (9a), wherein the area between these bounds encapsulates both exact and inexact loss values. Furthermore, as observed in Fig. 2, enhancing the accuracy of the lower-level solver leads to a narrower interval between the upper and lower bounds, thereby bringing the exact and inexact loss values closer.
4.2 Multinomial logistic loss
In order to compare MAID with HOAG [33] as a well-known gradient-based algorithm with inexact hypergradients, we replicate the multinomial logistic regression bilevel problem used in [33, 26] on the MNIST [^4] http://yann.lecun.com/exdb/mnist/ dataset. Similar to [33], we take the entire $M=60000$ training samples of MNIST and rescale them to $12\times 12$ pixels. Moreover, we utilize all $N=10000$ samples of the validation set for the upper-level problem. The bilevel problem corresponding to this model has the following form:
$\min_{\theta\in\mathbb{R}^{d}}f(\theta)$ $\coloneqq\sum_{i=1}^{N}\Psi(\bar{b}_{i},\bar{a}_{i}^{T}\hat{x}(\theta)),$ (23a)
$s.t.\quad\hat{x}(\theta)$ $\coloneqq\arg\min_{x\in\mathbb{R}^{p\times q}}\sum_{j=1}^{M}\Psi(b_{j},a_{j}^{T}x)+\frac{1}{2}\sum_{k=1}^{p}\sum_{l=1}^{q}e^{\theta[{(k-1)q+l}]}x_{k,l}^{2},$ (23b)
where $p=12\times 12=144$ represents the number of features, $q=10$ represents the number of classes, and $d=p\times q=1440$ denotes the number of parameters. $\Psi$ represents the multinomial logistic loss [29, Section 10.3], while $a_{j}\in\mathbb{R}^{p}$ and $b_{j}\in\mathbb{R}^{q}$ correspond to the training images and labels for each $1\leq j\leq M$. Similarly, $\bar{a}_{i}\in\mathbb{R}^{p}$ and $\bar{b}_{i}\in\mathbb{R}^{q}$ denote the validation images and labels in the upper-level problem for each $1\leq i\leq N$. To solve this problem, we employ MAID, alongside HOAG using its heuristic adaptive step size with default parameters. We set the initial accuracy for both algorithms to $\epsilon_{0}=\delta_{0}=10^{-1}$, as specified in [33]. Moreover, to ensure a fair comparison, similar to HOAG, we set the hyperparameters $\overline{\rho}=\frac{10}{9}$, $\underline{\rho}=0.5$, $\overline{\nu}=1.05$, and $\underline{\nu}=0.5$. In HOAG, the initial step size $\beta_{0}=\sqrt{d}/\|z_{0}\|$ was determined using the same strategy as outlined in [33]. We set the initial upper-level step size $\alpha_{0}$ in MAID equal to the initial step size in HOAG. Additionally, we consider a fixed total lower-level budget of $6\times 10^{5}$ iterations and a maximum of $300$ upper-level iterations.
Figure 3: The value of $\epsilon$, and upper-level loss per total lower-level computational cost in the algorithms MAID and HOAG for solving the problem (4.2). It illustrates that MAID adaptively adjusts the upper-level step size and the required accuracy for the solving the lower-level problem, while HOAG with the adaptive step size does not adjust the step size after a few iterations and exhibits a decreasing sequence for the accuracy $\epsilon$. In all HOAG and MAID settings, $\epsilon_{0}=\delta_{0}$ is set.
Considering the fixed computational budget, Fig. 3 offers a comparison between HOAG and MAID in terms of upper-level step size, the accuracy of the lower-level solver, and loss. Throughout this experiment, all configurations of HOAG, employing the adaptive step size strategy from [33] and depending on their accuracy regime, demonstrated the ability to effectively adjust the step size and accuracy, leading to loss reduction in (4.2) within a few upper-level iterations and their corresponding lower-level computational cost. However, HOAG’s adaptive step size eventually stops changing and becomes fixed at different levels, depending on the accuracy regime, while diminishing accuracy for the lower-level solver to negligible values. As a result, the lack of adaptivity in accuracy and reaching high accuracies lead to slower loss reduction using the quadratic and cubic sequences of accuracies.
In contrast, while using the same initial step size and accuracy as HOAG, MAID consistently identified an appropriate range for accuracy and upper-level step size. This led to driving the loss to lower values than those achieved by HOAG within the same budget. Moreover, upon fitting a polynomial regime to the accuracy sequence taken by MAID, the sequence $\epsilon_{k}={0.35}{(k+0.14)^{-1.05}}\epsilon_{0}-0.0009$ for $k=1,2,\dots$ could potentially be a suitable option for HOAG when compared to the three other named sequences. Importantly, since this sequence is summable, the theoretical convergence of HOAG also holds. Nevertheless, knowing such a sequence a priori is not straightforward, and MAID finds it adaptively.
In summation, MAID demonstrates superior adaptability and performance compared to HOAG, achieving lower loss values within the same lower-level computational budget. MAID achieved this by adaptively adjusting accuracy and step size, and showcased an enhanced efficiency. Additionally, the step size selection in MAID comes with a theoretical convergence guarantee, whereas the strategy in HOAG is heuristic.
4.3 Variational image denosing
4.3.1 Total variation denoising
In this part, we consider the smoothed version of the image-denoising model ROF [36] as implemented in [18]. Here, a bilevel problem is formulated to learn the smoothing and regularization parameters of total variation (TV) regularization. In contrast to the first numerical experiment in this section, access to the optimal hyperparameters is not possible a priori.
Similar to [18], we apply this model to $2D$ images and the training set is $m=25$ images of the Kodak dataset [^5] https://r0k.us/graphics/kodak/ resized to $96\times 96$ pixels, converted to grayscale, and Gaussian noise with $\mathcal{N}(0,\sigma^{2})$ with $\sigma=0.1$ is added. The bilevel problem we examine is as follows:
$\min_{\theta}\frac{1}{m}\sum_{t=1}^{m}\frac{1}{2}\|\hat{x}_{t}(\theta)-x^{*}\|^{2}$ (24a)
$s.t.\quad\hat{x}_{t}(\theta)=\arg\min_{x}{\frac{1}{2}\|x-y_{t}\|^{2}}+e^{\theta[1]}{\sum_{i,j}\sqrt{|\nabla x_{i}|^{2}+|\nabla x_{j}|^{2}+(e^{\theta[2]})^{2}}},$ (24b)
Here, $\nabla x_{i}$ and $\nabla x_{j}$ represent the $i$-th and $j$-th elements of the forward differences of $x$ in the horizontal and vertical directions, respectively. It is important to note that this problem satisfies the assumption 2. To initiate the process, we set $\theta_{0}=(-5,-5)$ and utilize Algorithm 1 along with the FISTA variant of DFO-LS with dynamic accuracy [18] to solve the problem (4.3.1). The DFO-LS with dynamic accuracy is a derivative-free bilevel algorithm that employs FISTA for solving the lower-level problem with an accuracy determined dynamically from the trust region radius in the derivative-free optimization algorithm for solving least-squares [12] in the upper-level problem. This algorithm serves as an efficient approach for solving bilevel problems and in particular for the problem (4.3.1). However, it should be noted that, as discussed in [15], in contrast with gradient-based algorithms, it may not scale well with the number of hyperparameters due to the nature of the derivative-free algorithm used for the upper-level problem. To maintain consistency with the first numerical example, we set a limit on the total number of lower-level iterations in this case.
Figure 4: The values of $\epsilon$, and upper-level loss per total lower-level computational cost in the algorithms MAID and DFO-LS for solving the problem (4.3.1). In various MAID settings, $\epsilon_{0}=\delta_{0}$ is set.
Illustrated in Fig. 3(a), MAID demonstrates that, regardless of the starting accuracy of the lower-level solver $\epsilon$—chosen with different orders of magnitude—it adapts the accuracy and converges to the same range after a few upper-level iterations. This highlights the robustness of MAID in the selection of an initial accuracy, enabling it to find values as large as necessary and as small as needed to make progress in backtracking line search. On the other hand, the FISTA variant of DFO-LS initially selects an equal or higher value of $\epsilon$ compared to MAID when using the same starting accuracies. However, at a certain point, it begins to significantly decrease $\epsilon$.
In terms of loss, as indicated in Fig. 3(b), with an equal total lower-level computational budget (assuming CG iterations of DFO-LS as zero), MAID surpasses the performance of the FISTA variant of DFO-LS, even in the case where the initial $\epsilon$ is chosen too small. All different setups of MAID converged to the loss value of $8.49$, while DFO-LS with $\epsilon_{0}=10^{-1}$ and $\epsilon_{0}=10^{-3}$ reached $8.97$ and $9.00$, respectively. Furthermore, it is important to note that DFO-LS does not utilize any hypergradient information, which could place it at a disadvantage when compared to gradient-based methods.
Now, to evaluate the effectiveness of the learned parameters obtained through solving (4.3.1) using MAID and dynamic DFO-LS, We solved the training problem (4.3.1) for $20$ images of the Kodak dataset, each of size $256\times 256$ with Gaussian noise $\sigma=0.1$, using MAID with $\epsilon_{0}=\delta_{0}=10^{-1}$ and the FISTA variant of dynamic DFO-LS with $\epsilon_{0}=\Delta_{0}=10^{-1}$, where $\Delta_{0}$ represents the trust region radius in the first upper-level iteration. Subsequently, we applied these learned parameters to the lower-level total variation denoising problem for a test noisy image of size $256\times 256$ and standard deviation $\sigma=0.1$.
Figure 5: Total variation denoising with learned parameters by MAID ($\epsilon_{0}=\delta_{0}=10^{-1}$) and dynamic DFO-LS ($\epsilon_{0}=10^{-1}$).
Figure 5 shows the denoising results for both algorithms. It displays the original and noisy image as well as the denoised images obtained using MAID and DFO-LS, respectively. Each denoised image is accompanied by its corresponding peak signal-to-noise ratio (PSNR) value. The learned hyperparameters using MAID result in higher PSNR values than DFO-LS learned hyperparameters on test images, which is consistent with the training results.
4.3.2 Field of experts denoising
In this part, we explore image denoising using a more data-adaptive regularizer known as the Field of Experts (FoE) approach. This approach was previously employed in a bilevel setting in [14] to learn the parameters of the regularizer in the variational setting. However, their choice of regularizer was non-convex and does not align with our assumptions. Therefore, we adopt the convex case, as described in [15], resulting in the following bilevel problem:
$\min_{\theta}\frac{1}{m}\sum_{i=1}^{m}\frac{1}{2}\|\hat{x}_{i}(\theta)-x^{*}\|^{2}$ (25a)
$s.t.\quad\hat{x}_{i}(\theta)=\arg\min_{x}\left\{\frac{1}{2}\|x-y_{i}\|^{2}+e^{\theta[0]}\sum_{k=1}^{K}e^{\theta[k]}\|c_{k}\ast x\|_{\theta[{K+k}]}\right\},\quad i=1,\dots,m,$ (25b)
where each $c_{k}$ is a $2D$ convolution kernel. We set $K=30$ and take each $c_{k}$ as a $7\times 7$ kernel with elements $[\theta[{2K+49(k-1)+1}],\dots,\theta[{2K+49k}]]$. Also, $\|x\|_{\upsilon}=\sum_{j=1}^{n}\sqrt{x_{j}^{2}+\upsilon^{2}}$ represents the smoothed $1$-norm with smoothing parameter $\upsilon$ for any $x\in\mathbb{R}^{n}$. So, this problem has $1531$ hyperparameters that we represent by $\theta$. We initialize the kernels by drawing randomly from a Gaussian distribution with mean 00 and standard deviation $1$, and we set the initial weights to $e^{-3}$. During training, similar to TV denoising, we use 25 images from the Kodak dataset and rescale them to $96\times 96$ pixels, converting them to grayscale. Additionally, we add Gaussian noise with $\mathcal{N}(0,\sigma^{2})$, where $\sigma=0.1$, to the ground truth images.
Figure 6: Investigating the impact of MAID configurations on the accuracy of the lower-level solver, and loss at each lower-level computation unit with varied $\epsilon_{0}=\delta_{0}$ for solving the problem (4.3.2).
As seen in Fig. 5(a), by dynamically adjusting the accuracy within the framework of MAID, After a few upper-level iterations in various settings, the algorithm ultimately selects the same range of accuracies. Akin to the patterns observed in previous numerical experiments, choosing a too-large initial accuracy $\epsilon_{0}=10$ in this case prompts the algorithm to undergo reductions and recalculations to find a descent direction, resulting in higher computational costs initially. However, after sufficiently reducing $\epsilon$, it manages to make progress with a reasonable computational cost.
The advantage of adaptive accuracy in MAID becomes evident when comparing it to fixed accuracies. Using a low fixed accuracy $\epsilon_{0}=10^{-1}$ along with the line search mechanism of MAID, once can see a rapid progress initially, as indicated by the pink dashed curve, but it reaches a plateau after a few upper-level iterations. Furthermore, it does not use the full lower-level budget because the upper-level progress stalls at some point, and due to warm-starting, the lower-level solution is accurate enough and does not require many more iterations. A medium fixed accuracy, such as $10^{-3}$, performs comparably to adaptive cases; however, it cannot be known a priori and must be chosen experimentally. A high fixed accuracy, like $\epsilon_{0}=10^{-5}$, appears sufficiently small but exhausts the lower-level solver and runs out of budget while achieving higher loss. All adaptive instances demonstrate enhanced overall efficiency compared to their fixed counterparts, as showcased in Fig. 5(b).
Noting the importance of the problem (4.3.2) as a large-scale and data-adaptive problem, to evaluate the learned parameters, we utilized the same set of $20$ images used in the total variation problem, each of size $256\times 256$. Additionally, we employed $48$ filters of size $7\times 7$, initialized with zero-mean random filters. After training using MAID with $\epsilon_{0}=\delta_{0}=5\times 10^{-1}$, we evaluated the learned model on a noisy test image of size $256\times 256$ pixels with Gaussian noise of $\sigma=0.1$, as depicted in Fig. 6(a) and Fig. 6(b) respectively.
The quality of the learned parameters and, consequently, the denoised image obtained through MAID, are illustrated in Fig. 6(c). The denoised image by the learned FoE model achieved a PSNR of $30.01$ dB, which is higher than the PSNR of $29.33$ dB obtained by TV with learned parameters on the same data.
Figure 7: Field of experts denoising with learned parameters utilizing MAID.
5 Conclusion
MAID is the first algorithm which actively exploits the fundamental trade-off in the accuracy and computational cost in estimating function values and gradients in bilevel learning problems. Rather than a-priori selected, MAID selects accuracies for both function values and gradients following the overarching paradigm to be as low accurate as possible to save computational cost but as accurate as needed to guarantee global convergence to critical points. A key novelty is that MAID performs backtracking using inexact computations to assure monotonic descent. Our numerical results underscore the superiority of MAID over state-of-the-art derivative-free and first-order methods for bilevel learning across a range of problems in data science. Importantly, it demonstrates remarkable robustness in the face of hyperparameters such as the initial accuracy and starting step size choices, alleviating the need for tuning these algorithm parameters specifically for each application.
Acknowledgments
This research made use of Hex, the GPU Cloud in the Department of Computer Science at the University of Bath.